How many litres each of 60% ammonia and 80% ammonia must be mixed to obtain 20 litres of 75% ammonia ?
Answer by fractalier(6550)

Call the amount of 60% ammonia, x.
Then the amount of 80% ammonia will be 20 - x.
The set up is
.60x + .80(20-x) = .75(20)
.60x + 16 - .80x = 15
-.20x = -1
x = 5 liters at 60% and then
20-x = 15 liters at 80%